Prev Next

SW&IT Meeting Minutes 2019-10-14

Attendees: Andrew, David, Igor, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20191007
LR20191007 - Bec, Matthieu 80.0
LR20191007 - Holmberg, Carl 40.0
LR20191007 - Colson, Andrew 40.0
LR20191007 - Edmeades, David 29.0
LR20191007 - Sola, Igor 0.0
LR20191007 - Kubanek, Petr 0.0
LR20191007 - Hooper, Stephen 80.0
LR20191007 - Zhang, Yang 40.0

News from management

  • Dec 4 - Jan 12 CV and north, move to 3rd floor in December / January - needs for network
    • 2nd part: Jan 14 - 21 (tentative), open space location - details will be further discussed
  • Shutdown procedure and lessons learnt. Review planned shutdown, emergency shutdown, etc.

Previous Action Items

Open topics

  • do we have adsec/wfs spares?
  • new AO/4D reduction VM server, AO data plan
  • PEPSI cluster
  • catalogs.lbto.org
  • iLocater 10Gb, snmp
  • FosWIKI deaths - reinstall new OS?

Team reports

DE

  • Projector replacement
  • PEPSI sketch sent, team less ready than we thought, hard to tell if they know what they need
  • Search for storage (timeline?) 300TiB to start?
    • Dell/EMC options
      • Isilon
      • Unity
      • Have to go through VAR, is painful. Process started
      • Will make comparison document/presentation when I have current numbers
    • Dell option
      • SCv3000
      • Requires server hosts to accept FC and output NFS/CIFS
  • VMWare fixed quote RSN--use old quote for PEPSI?
  • VMWare conversation re: PEPSI Monday morning
    • Should be covered under second 3-host license
  • First SSD failure in VM host
    • Alarm passed to VMWare
    • Rebuild status needs additional software that requires a reboot to install.
  • Will poke Netops re: VRF extension to mountain
  • Revisit new KVM: https://150.135.245.155 swg/b1nocular

SH

  • Wiki re-install, reconfigure
  • IRAF issues on rm507 computers -- fixed, but not sure what happened to break it.
  • Ricoh printer issues for Windows' users.
  • Weather serial reader -- first pass
  • ARGOS remote desktop -- not a great idea, but should we care? Argos-SX works from new rm507 computers, x2go core-dumps on older computers.

next:

  • weather serial testing
  • rm507 -- new observer script
  • Italian VPN and Polycom
  • Fedora 31 beta out -- RPM builds.

AC

  • ARGOS
    • supported daytime and eng tests
    • Ran ARGOS and ESM in parallel
    • changed SX AdSec configuration from SOUL to UAO
    • x2go session for argos-sx/dx cofigured in rm-507 (SH)
      • Need to reconfigure navigation bar menu options for various applications
  • On-call support
    • DIMM not "seeing" because of lost connection to camera. Solution: restart DIMM electronics. Instructions to be added to wiki.
  • IT7882 - FacSum GCS images
    • WFS and guider images were not displaying because fitstopnm needed to be installed and the images were "old." New images are downloaded and shown only if more than 60 seconds have elapsed since prior image. Closed.

next
  • ARGOS
    • revision control and RPM
    • 6-head/4k monitor setup
  • OAC-GCS interface - catch up with PK

YZ

  • AO-Cravat
    • Started major rework of the web application.
    • Made lots of progress on plotting (using plotly.js).
    • Next: continue to improve plotting and starting to rework Steve's Strehl code.
  • LBC/MODS
    • Finished LBC telemetry integration with web-dms system.
    • A major work for the next few weeks would be to rewrite the LBC non-sidereal guiding algorithm (to address IT#7847).
    • Need to address two other MODS related issues (IT#6939 and IT#7855).

MB

next
  • DMS
    • alta/repository final touch
    • https?
    • refactor, git cleanup, release v1.0
  • PEPSI ctd.
  • catch up with PK (tcs-snoop), SH (obswkup), XZ (AO data)

IS

Last week:
  • OT
    • LUCI ARC off lamps
    • Look at DEC column in JSKY catalogs. It displays the wrong column.
  • PIT
    • backend messages were not propagating to GUI

  • LUCI
    • Looking for metrics on the contention inside LUCI java daemons, specifically the Journalizer.
      • jmc, jvisualvm, jcmd

This week
  • LUCI
    • Keep looking at Journalizer performance and possibly break journalizer into several specialized intances: * A Journalizer GUI * A FITS generation Journalizer?
  • OT
    • FITS loading in PE issues
    • Tooltips for hi/low value cutoffs still not correct in sequence.

CH

Last week:
  • CHAOS:
    • Collected Zabbix data collection features, web dashboard customization.
  • SupportScripts:
    • Attempted to package pyds9, not quite there.

This week:
  • CHAOS: Finish Zabbix paper.
  • SupportScripts: Finish packaging pyds9, pacakge ds9.
Topic revision: r17 - 15 Oct 2019, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback